I can't for the life of me get to the right answer here. Can you help?
2A/3=8+4A, solve for A

Answer by checkley77(12569)

2A/3=8+4A
2A=3(8+4A)
2A=24+12A
2A-12A=24
-10A=24
A=24/-2
A=-2.4 ANSWER.
PROOF:
2*-2.4/3=8+4*-2.4
-4.8/3=8-9.6
-1.6=-1.6